OpenStudy (cyclone3260):

Help with Algebra 2 please? 1.(x+y)^3 I got x^3 + 6x^2 +12x+8 Am I correct?? 2. (x+y)^4 3. (x+1)^4 4. (x - 3y)^3 Help please ill try to give as MUCH points as I can I promise

OpenStudy (anonymous):

Have you tried. https://www.wolframalpha.com/

OpenStudy (jdoe0001):

\(\bf (x+y)^3\implies (x+y)(x+y)(x+y) \ne x^3 + 6x^2 +12x+8 \)

OpenStudy (cwrw238):

do it in 2 steps (x + 3)(x + 3) distribute x and then +3 over the second bracket = x^2 + 3x + 3x + 9 = x^2 + 6x + 9 now work out (x + 3)(x^2 + 6x + 9) by a similar method

OpenStudy (cyclone3260):

Thanks! :) What about the rest though?

OpenStudy (cwrw238):

they can all be done in much the same way or you can use the Binomial theorem - have you done that yet?